OpenStudy (anonymous):

Find f(5) for this sequence: f(1)=2 and f(2)=3, f(n)= f(n-1)+f(n-2), for n>2

OpenStudy (anonymous):

\[f(1)=2,f(2)=3,f(3)=f(2)+f(1)=2+3=5\]

OpenStudy (anonymous):

and so \(f(4)=f(3)+f(2)=3+5=8\)

OpenStudy (anonymous):

one to go

OpenStudy (anonymous):

f(5)=f(4)+f(3) 8+5 = 13 Can You tell me if is ok please